### v1

# Static Word Embeddings for Sentence Semantic Representation

Abstract
We propose new static word embeddings optimised for sentence semantic representation. We first extract word embeddings from a pre-trained Sentence Transformer, and improve them with sentence-level principal component analysis, followed by either knowledge distillation or contrastive learning. During inference, we represent sentences by simply averaging word embeddings, which requires little computational cost. We evaluate models on both monolingual and cross-lingual tasks and show that our model substantially outperforms existing static models on sentence semantic tasks, and even rivals a basic Sentence Transformer model (SimCSE) on some data sets. Lastly, we perform a variety of analyses and show that our method successfully removes word embedding components that are irrelevant to sentence semantics, and adjusts the vector norms based on the influence of words on sentence semantics.
Static Word Embeddings for Sentence Semantic Representation
Takashi Wada   Yuki Hirakawa   Ryotaro Shimizu   Takahiro Kawashima   Yuki Saito ZOZO Research {firstname.lastname}@zozo.com
1 Introduction
In natural language processing (NLP), it has been a major topic to learn fixed-length embeddings that represent sentence semantics. To achieve this goal, most existing methods fine-tune pre-trained language models (LMs) such as BERT Devlin et al. (2019) using labelled data of various NLP tasks (e.g. natural language inference). Recent work further employs large language models (LLMs) such as LLaMA 3 AI@Meta (2024) and shows that scaling up models can yield better performance BehnamGhader et al. (2024). However, this approach requires a large amount of computational cost, making it difficult to process billions of sentences cost-efficiently or deploy models on resource-constrained devices such as smartphones. As such, it is crucial to explore more cost-efficient models that work well without GPUs, and yet this direction remains largely overlooked in the research community.
In this work, we propose new static word embeddings (SWEs) — both monolingual and cross-lingual ones — that can represent sentence semantics fairly well with little computational cost. Specifically, our method first extracts SWEs from a pre-trained Sentence Transformer Vaswani et al. (2017); Reimers and Gurevych (2019), and improves them with sentence-level principal component analysis (PCA) followed by either knowledge distillation or contrastive learning. During inference, we encode sentences by simply looking up SWEs and taking their average, which can be done comfortably with limited resources. Our main contributions are: (1) we propose new SWEs that substantially outperform existing SWEs on a semantic textual similarity (STS) task, and even surpass SimCSE Gao et al. (2021) on Massive Text Embedding Benchmark Muennighoff et al. (2023) in English; (2) we propose cross-lingual SWEs optimised for sentence representation, which achieve surprisingly good performance on translation retrieval even between distant languages (e.g. 94.6% in F1 scores on English–Chinese); and (3) we conduct a variety of analyses and provide new insights into how to obtain SWEs suitable for sentence representation. Our SWEs will be available at https://github.com/twadada/swe4semantics.
2 Related Work
Most recent text embedding models are built on pre-trained Transformer LMs. For instance, GTE Li et al. (2023) fine-tunes BERT with contrastive learning on labelled data from various sources (e.g. STS data sets). Later, it is extended to the multilingual model called mGTE Zhang et al. (2024), which is trained on both monolingual and cross-lingual data sets. LLM2Vec BehnamGhader et al. (2024) scales up the backbone model and fine-tunes LLMs such as LLaMA with contrastive learning. Similarly, NV-Embed Lee et al. (2025) fine-tunes Mistral-7B Jiang et al. (2023) with contrastive learning and also trains an additional attention layer on top. While these models achieve impressive results, they are computationally expensive and inefficient for real-world applications. To mitigate this, several methods to train compact models are proposed such as knowledge distillation Hinton et al. (2015); Xu et al. (2023), but they are mostly applied to Transformer models, which are designed with GPU acceleration in mind.
On the other hand, little attention has been paid to building cost-efficient models (esp. cross-lingual ones) that can be easily run on CPUs. Although such models inevitably lag behind heavy-weight models in accuracy, they offer significant advantages in cost and runtime efficiency, and can be applied to, for instance, quickly mining semantically similar content from billions of posts on social media. One traditional model in this direction is Sent2Vec Pagliardini et al. (2018), which extends the CBOW algorithm Mikolov et al. (2013a) and represents sentences by averaging static word and bigram embeddings. Very recently, two static word embedding models called WordLlama Miller (2024) and Model2Vec Tulkens and van Dongen (2024) have been released from different GitHub projects (MIT license), [^1] https://github.com/dleemiller/WordLlama and https://github.com/MinishLab/model2vec, respectively. and demonstrate strong performance on sentence-level tasks in English. Crucially, however, they do not publish a paper or detailed technical report on these models, leaving their tuning processes unclear and lacking the analyses of their embeddings (unlike our work). Based on their code and documentations, both models extract SWEs from Transformer models. Specifically, WordLlama extracts SWEs from the input layer of LLMs like LLaMa 3 AI@Meta (2024), and refines them with contrastive learning on various labelled data such as STS and text retrieval data sets. Model2Vec extracts SWEs from a Transformer text encoder called BGE-base Xiao et al. (2023), by feeding each word $w$ in a vocabulary into BGE without context (i.e. “[CLS], $w$, [SEP]”) and using the output as the embedding of $w$. It then fine-tunes the SWEs by minimising the mean cosine distance between the sentence embeddings produced by BGE and the average of the SWEs for the same input text. To reduce the dimensionality, both WordLlama and Model2Vec perform word-level PCA as a post-processing step, and Model2Vec also applies the smooth inverse frequency (SIF) heuristic Arora et al. (2017), which multiplies the norm of $w$ by $\tfrac{\alpha}{\alpha+\mathrm{p}(w)}$, where $\mathrm{p}(w)$ denotes the uni-gram probability of $w$ and $\alpha=0.001$.
Amid the rapid growth in English models, there is a paucity of cross-lingual SWEs optimised for sentence representation, as most existing ones are designed for word-level tasks such as bilingual lexicon induction and cross-lingual word alignment Mikolov et al. (2013b); Xing et al. (2015); Conneau et al. (2018); Artetxe et al. (2018); Wada et al. (2019, 2021). Although Model2Vec also releases a multilingual model that extracts SWEs from LaBSE Feng et al. (2022), its cross-lingual performance is limited as we show in our experiments. We address this problem and propose effective SWEs not only for monolingual tasks but also for cross-lingual ones, such as (sentence) translation retrieval.
3 Methodology
Our proposed method obtains SWEs in three steps: (1) embedding extraction from Sentence Transformer (Sec. 3.1); (2) dimensionality reduction with sentence-level PCA (Sec. 3.2); and (3) embedding refinement with knowledge distillation or contrastive learning (Sec. 3.3). Only the last step involves the training (fine-tuning) of SWEs.
3.1 Embedding Extraction
We first extract SWEs from a pre-trained Sentence Transformer (ST). Specifically, to obtain the embedding of a word $w\in V$ (denoted as $\mathrm{E}(w)$), we sample $N=100$ sentences $z^{(w)}_{i}~{}(i=1,2,...,N)$ that contain $w$ from a large corpus. Then, we encode each sentence $z^{(w)}_{i}$ using ST, which produces $|z^{(w)}_{i}|$ vectors at the last Transformer layer ($|z^{(w)}_{i}|$ denotes the number of tokens in $z^{(w)}_{i}$). We then retrieve the vector at the position of $w$ (denoted as $\mathrm{ST}(w,z^{(w)}_{i})$), and obtain $\mathrm{E}(w)$ as:
$\mathrm{E}(w)=\frac{1}{N}\sum_{i=1}^{N}{\mathrm{ST}(w,z^{(w)}_{i})}.$ (1)
This way, we decontextualise the embeddings and obtain the SWE of $w$. Furthermore, we can also generate cross-lingual SWEs using multilingual ST, as it embeds different languages in the same space. For the vocabulary $V$, we include more words than the ST’s original vocabulary, and if $w$ is tokenised into subwords, we average its subword embeddings to obtain $\mathrm{ST}(w,z^{(w)}_{i})$ in Eqn. (1).
The idea of extracting SWEs from Transformers has been explored in previous studies Ethayarajh (2019); Bommasani et al. (2020); Wada et al. (2022), which show that SWEs extracted from masked LMs perform well on word-level tasks (e.g. lexical substitution). However, we find that these embeddings are not effective for sentence-level tasks, and hence we refine them in the following steps.
3.2 Sentence-level PCA
Next, we apply sentence-level PCA to the extracted SWEs in Section 3.1 to improve performance and reduce the embedding dimensionality. In previous work (including WordLlama and Model2Vec), the dimensionality reduction of SWEs is typically done by applying PCA on a word embedding matrix $Y\in\mathbb{R}^{|V|\times d}$, where $d$ denotes the embedding dimension; we denote this as word-level PCA. On the other hand, we apply PCA to a sentence embedding matrix (as explained below), which emphasises the variance of sentences rather than words and improves performance on sentence-level tasks.
Specifically, we randomly sample $M=100\mathrm{k}$ sentences from those used in Section 3.1, and represent each sentence by averaging SWEs. When training cross-lingual SWEs, we sample $M$ sentences from each language and concatenate them. Given the sentence embedding matrix $X\in\mathbb{R}^{M\times d}$ (or $\in\mathbb{R}^{2M\times d}$) and its mean vector $\bar{X}\in\mathbb{R}^{d}$, we apply PCA to the centred matrix $X-\bar{X}$ and obtain the transformation matrix $W$. This matrix transforms the embedding space onto new axes that effectively capture the variance of the sentence representations. Importantly, since we represent each sentence by averaging SWEs, we can pre-transform the SWE $\mathrm{E}(w)$ into $\mathrm{\hat{E}}(w)=W^{\top}(\mathrm{E}(w)-\bar{X})$ for each word $w\in V$ once and save $\mathrm{\hat{E}}(w)$ in memory, as $W^{\top}(\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{E}(w^{\prime})-\bar{X})=\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{\hat{E}}(w^{\prime})$.
To reduce the dimensionality using PCA, a common approach is to keep the principal components (PCs) with the $d^{\prime}(<d)$ largest eigenvalues. However, we instead discard the first $r$ PCs, and keep the $(r+1)$-th to $(r+d^{\prime})$-th components. This is based on the previous finding that removing a few dominant PCs from SWEs can counter-intuitively improve the embedding quality Mu and Viswanath (2018), and we actually get positive results in our experiments especially in cross-lingual tasks (e.g. +49.4% in F1 scores). Intriguingly, our analysis (Sec. 5.2) reveals that it helps remove information not relevant to sentence semantics. Following Mu and Viswanath (2018), we set $r$ to $\lfloor\frac{d}{100}\rfloor$, and refer to this method as ABTT (All-But-The-Top).
3.3 Embedding Refinement
Lastly, we refine the dimensionality-reduced embeddings after PCA. For monolingual SWEs, we perform teacher-student knowledge distillation, where we fine-tune our SWEs (= student) to reproduce the sentence similarities calculated by ST (= teacher), which is freezed during distillation; hence, it is not computationally expensive and can be done without labelled data (unlike WordLlama). Formally, for a mini-batch of $K=128$ sentences, we minimise the following loss $\mathcal{L}$:
$\mathcal{L}=-\frac{1}{K}\sum_{i=1}^{K}\sum_{j\neq i}^{K}\sigma\!\left({\mathbf{T},i,j}\right)\log\sigma\!\left({\mathbf{S},i,j}\right),$ (2)
$\sigma(\mathbf{X},i,j)=\frac{e^{\frac{x_{i,j}}{\tau}}}{\sum_{k\neq i}^{K}e^{\frac{x_{i,k}}{\tau}}},$ (3)
where $\mathbf{S},\mathbf{T}\in\mathbb{R}^{K\times K}$ denote the similarity matrices calculated by the teacher and student models, with the value at the $i$-th row and $j$-th column corresponding to the cosine similarity between the $i$-th and $j$-th sentences in the mini-batch. The temperature $\tau$ is set to $0.05$ as commonly done in contrastive learning Gao et al. (2021). Note that the student model represents each sentence by averaging SWEs obtained in Section 3.2. Since the teacher and student models calculate $S$ and $T$ independently, their embedding spaces can be different and that makes it possible to fine-tune the dimensionality-reduced embeddings after PCA. In contrast, Model2Vec minimises the cosine distance of embeddings output by ST and SWEs, requiring both models to share the same embedding space; hence, it applies (word-level) PCA after fine-tuning, which can diminish the benefits of fine-tuning and hence relies on the SIF heuristic to improve performance. Section 5.3 shows that our method yields a similar effect to SIF, but in a more refined manner.
For cross-lingual SWEs, we find that performing knowledge distillation for each language leads to worse results on cross-lingual tasks, as it does not encourage SWEs to be mapped into the same cross-lingual space. Therefore, we instead make use of a parallel corpus as a source of knowledge, and fine-tune the cross-lingual SWEs in languages $\ell 1$ and $\ell 2$ by minimising the following loss $\mathcal{L}_{CL}$:
$\mathcal{L}_{CL}=-\frac{1}{K}\sum_{i=1}^{K}(\log\sigma^{\prime}(\mathbf{U},i)+\log\sigma^{\prime}(\mathbf{U}^{\top},i)),$ (4)
$\sigma^{\prime}(\mathbf{X},i)=\frac{e^{\frac{x_{i,i}}{\tau}}}{\sum_{k=1}^{K}e^{\frac{x_{i,k}}{\tau}}},$ (5)
where $\mathbf{U}\in\mathbb{R}^{K\times K}$ denotes the similarity matrix calculated by SWEs, and $u_{i,j}$ denotes the cosine similarity between the $i$-th sentence in $\ell 1$ and $j$-th sentence in $\ell 2$ in the mini-batch. The $i$-th sentences in $\ell 1$ and $\ell 2$ are a pair of translations, and we create mini-batches by sampling $K$ translation pairs from a parallel corpus. [^2] We transpose $U$ in Eqn. (4) to calculate the denominator in Eqn. (5) for each language.
4 Experiments
4.1 Our Models
We train our monolingual (English) SWEs using GTE-base Li et al. (2023), which achieves strong performance on Massive Text Embedding Benchmark (MTEB) among BERT-sized models. We sample sentences from a large-scale unlabelled corpus (CC-100 Wenzek et al. (2020); Conneau et al. (2020)) and use them in each step of our method. For the vocabulary $V$, we include the 150k most frequent words (case-insensitive) in CC-100. To generate cross-lingual SWEs, we use mGTE Zhang et al. (2024), the multilingual version of GTE trained on 75 languages. We train cross-lingual embeddings for three language pairs, namely English–German (en–de), English–Chinese (en–zh), and English–Japanese (en–ja). We sample short sentences from large-scale parallel corpora (CC-Matrix Schwenk et al. (2021)) and use them to extract SWEs in Section 3.1 and calculate the loss in Eqn. (4). For en–ja, we also use JParaCrawl v3.0 Morishita et al. (2020, 2022) to augment data. For $V$ in English and German, we choose the 150k most frequent words (case-sensitive) in the corresponding corpora, and 30k most frequent tokens (which can be a subword or consist of multiple words) in Japanese and Chinese. [^3] This is because mGTE’s tokeniser segments sentences into tokens without performing word segmentation in both languages (which have no whitespace word boundary).
In Section 3.2, we reduce the dimensionality to $d^{\prime}\in\{256,512\}$, [^4] In most cases, 99% of the variance can be explained with $d^{\prime}=512$ when the original dimensionality $d$ is $768$. and in Section 3.3 we fine-tune SWEs using Adam Kingma and Ba (2015) with the learning rate of 0.001 for 30k steps, and apply early stopping based on the validation loss; the train and validation data consist of sentences sampled from those used in Section 3.1. During inference, we tokenise the text $z$ into words (or tokens in Japanese and Chinese) and strip punctuation, and average SWEs to represent $z$. When a word is not in $V$, we tokenise it into subwords using (m)GTE’s tokeniser, and truncate it by removing the last subword until the remaining string is found in $V$ [^5] For instance, if tokeniser is out of vocabulary, we split it into subwords (e.g. token #ise #r), and strip the last subword(s). Then, if tokenise (or token) is in $V$, we use its embedding.
Models ($d$) STS12 STS13 STS14 STS15 STS16 STS17 STS22 STS-B SICK-R BIO Sentence Transformers (STs) MiniLM-L6 (384) 72.4 80.6 75.6 85.4 79.0 87.6 67.2 82.0 77.6 81.6 SimCSE (768) 75.3 84.7 80.2 85.4 80.8 89.4 62.0 84.2 80.8 68.4 BGE-base (768) 78.0 84.2 82.3 88.0 85.5 86.4 65.9 86.4 80.3 86.9 GTE-base (768) 74.4 84.7 80.1 87.2 85.0 90.6 68.6 86.0 79.4 83.6 LLM2Vec (4,096) 79.3 84.8 82.9 88.1 86.5 89.6 67.7 88.0 83.9 84.9 Static Word Embeddings (SWEs) fastText (300) 57.2 69.2 62.8 73.0 64.2 70.2 52.1 56.5 60.1 63.0 Sent2Vec (700) 54.1 66.3 65.8 78.0 70.6 82.2 54.5 68.0 63.2 55.2 WordLlama (512) 63.9 73.3 69.1 81.5 76.1 85.2 60.2 77.0 67.0 71.3 Model2Vec (512) 62.7 77.6 72.9 80.8 76.9 87.1 64.3 76.8 65.7 77.6 OURS (256) 68.3 79.3 75.9 83.1 79.4 86.1 59.3 79.2 68.0 80.3
Table 1: Results on STS data sets. The best scores among the ST and SWE models are boldfaced.
4.2 Baselines
In monolingual experiments, we include two traditional SWEs as our baselines, namely fastText Bojanowski et al. (2017) and Sent2vec. Both models are self-supervised based on word co-occurrence information, and Sent2vec performs better on sentence-level tasks. We also compare our model against two recent GitHub-hosted models with no paper: WordLlama and Model2Vec. We use wordllama-l3-supercat for WordLlama, which is based on LLaMa 3, and potion-base-32M for Model2Vec, the latest and state-of-the-art model (released in January 2025) based on BGE-base. SWE models, including ours produce sentence vectors by averaging SWEs, followed by L2-normalisation.
In cross-lingual experiments, we compare our model against the multilingual Model2Vec model (M2V_multilingual_output). For English–German, we also evaluate cross-lingual SWEs available at the MUSE GitHub repository [^6] https://github.com/facebookresearch/MUSE Conneau et al. (2018), which align English and German fastText embeddings using a bilingual dictionary and have been widely used in the research community.
Models ($d$) $100\rho$ runtime (s) GPU CPU MiniLM-L6 (384) 85.4 1.7 8.8 GTE-base (768) 87.2 3.4 52.7 LLM2Vec (4,096) 88.1 30.3 $>$10,000 WordLlama (512) 81.5 0.4 Model2Vec (512) 80.8 0.3 OURS (256) 83.1 0.4
Table 2: Scores and runtime (in second) on STS15. The runtime is averaged over three independent runs, and is measured on the same virtual machine with or without enabling a GPU (NVIDIA A100-SXM4-40GB).
4.3 Experiments on English Tasks
We first evaluate monolingual (English) models on semantic textual similarity (STS), a well-established task for sentence semantic evaluation. It is a task of calculating textual similarity for each sentence pair, and models are evaluated based on Spearman’s rank correlation $\rho$ between their predictions and human judgements. Table 1 shows the results on STS12–17, 22, STS Benchmark (STS-B), SICK-R, and BIOSSES (BIO) data sets Agirre et al. (2012, 2013); Bandhakavi et al. (2014); Biçici (2015); Agirre et al. (2016); Chen et al. (2022); Marelli et al. (2014); Soğancıoğlu et al. (2017). The first five rows show performance of STs for reference (not as baselines of our model): SimCSE [^7] We use the supervised one, which performs better than the unsupervised one. is a pioneering ST, which often serves as a strong baseline in STS research; MiniLM-L6 is a popular light-weight ST based on MiniLM Wang et al. (2020); BGE and GTE are the models used to train Model2Vec and OURS; and LLM2Vec is a recent large model built on LLaMa-3-8B. The table shows that OURS substantially outperforms all SME baselines on most data sets, even with the smallest dimensionality ($d=256$). Our model also outperforms MiniLM on STS14 and STS16, and SimCSE on BIOSSES, where the sentences are sampled from biomedical domains, indicating the robustness of OURS to various types of text. Table 2 compares the scores and runtime on STS15. While falling behind STs in accuracy, OURS (and other SWEs) can run extremely fast on CPUs (20 times faster than MiniLM) without sacrificing the accuracy very much.
Models ($d$) Avg-s2s Avg MiniLM-L6 (384) 65.95 56.26 SimCSE (768) 62.75 48.86 BGE-base (768) 71.18 63.55 GTE-base (768 71.51 64.11 LLM2Vec (4,096) 72.94 65.01 fastText (300) 53.68 43.05 Sent2Vec (700) 56.33 46.29 WordLlama (512) 60.40 50.23 Model2Vec (512) 62.37 51.33 OURS (256) 63.76 51.97 OURS (512) 64.06 52.35
Table 3: Results on MTEB experiments in English. The best scores among STs and SWEs are bold-faced. Avg of Model2Vec is slightly lower the reported one due to a small bug in their MS MARCO evaluation.
Additionally, to further evaluate the effectiveness of our model on other tasks such as sentence classification, we also run experiments on Massive Text Embedding Benchmark (MTEB) Muennighoff et al. (2023). It consists of 56 data sets covering 7 different tasks, namely classification, pair classification, clustering, reranking, retrieval, STS, and summarisation evaluation. On each task, a model encodes text and returns its fixed-length vector, which is then used on downstream tasks (e.g. for calculating textual features or similarities). Models are evaluated using task-specific metrics (e.g. precision) on each data set. Since MTEB contains both sentence-level and document-level tasks, and our focus is on sentence semantic representation, we report two scores: Avg, the average across all 56 data sets; and Avg-s2s, the average across the 33 data sets labelled as “s2s”, where the input text is a sentence rather than a document. Table 3 presents the results, showing that our model substantially outperforms the other SWEs in Avg and Avg-s2s, even with the smallest embedding size. Moreover, our model outperforms SimCSE in both Avg and Avg-s2s, which is surprising given that our model does not use any sequential function like Transformer. Interestingly, we also find that ensembling OURS with WordLlama and Model2Vec further boosts performance (64.55 and 53.28 in Avg-s2s and Avg), which we describe in Appendix B. We also show the scores on each task in Table 10 in Appendix.
Models ($d$) BUCC Tatoeba en-de en-zh en-de en-zh en-ja mGTE (768) 98.6 98.2 97.8 95.7 93.1 MUSE (300) 46.6 44.3 Model2Vec (256) 60.0 0.0 73.5 26.6 15.3 OURS (256) 96.3 94.6 95.1 87.9 79.0
Table 4: Results (F1) on translation retrieval tasks.
Models ($d$) STS17 STS22 en-de en-de en-zh mGTE (768) 84.7 62.3 72.9 MUSE (300) 11.2 36.9 Model2Vec (256) 54.2 39.3 26.2 OURS (256) 70.5 62.0 72.6
Table 5: Results on cross-lingual and monolingual (German and Chinese) STS tasks.
4.4 Experiments on Cross-Lingual Tasks
We first evaluate cross-lingual SWEs on translation retrieval, a task of retrieving the translation of the input text based on cosine similarity. Table 4 shows the results (F1 scores) on BUCC Zweigenbaum et al. (2017) and Tatoeba Artetxe and Schwenk (2019); community (2021) data sets for three language pairs: English–{German, Chinese, Japanese}. The first row shows the performance of mGTE, and the rest are cross-lingual SWEs. It clearly shows that OURS outperforms the SWE baselines by a large margin, and performs fairly well on all language pairs (e.g. 96.3% for English-German on BUCC). This suggests the applicability of our model as a light-weight method for extracting translation candidates, which could be further validated by computationally expensive models such as mGTE. Lastly, we also evaluate SWEs on cross-lingual STS (STS17 and STS22), a task of measuring textual similarity across two different languages, and Table 5 presents the results. Again, OURS is the best performing SWE, demonstrating its effectiveness on cross-lingual tasks.
en en-de en-zh en-ja Section 3.1 44.2 51.6 45.0 28.6 Section 3.2 49.9 93.8 88.6 74.6      w/o ABTT 48.8 49.6 39.2 33.2 Section 3.3 52.0 95.7 91.2 79.0
Table 6: Performance of our SWEs obtained at each step. The “en” column denotes Avg on MTEB, and the rest are the average scores on BUCC and Tatoeba. The third row shows the results without ABTT in Section 3.2.
SimCSE Nomic GTE $d$ 768 768 384 768 1,024 ST 62.8 69.5 69.3 71.5 71.7 OURS 61.2 62.7 60.8 63.8 62.9
Table 7: Performance (Avg-s2s on MTEB) of different STs and OURS (256) trained with them.
5 Analysis
Figure 1: Scatter plots of our cross-lingual SWEs in English (blue) and Japanese (orange). The leftmost shows the 1st and 2nd principal components, and the rest are t-SNE visualisation before and after applying PCA wo/w ABTT.
5.1 Ablation Studies
We evaluate our SWEs ($d=256$) obtained in each section and Table 6 shows the results. The column under “en” shows performance on MTEB (Avg), indicating that sentence-level PCA and knowledge distillation substantially improve the SWEs extracted in Section 3.1. Other important findings that are omitted from Table 6 include: (1) applying word-level PCA in Section 3.2 performs worse than sentence-level PCA (48.1 vs. 49.9); and (2) performing knowledge distillation without PCA brings no improvement (44.4) [^8] Performance increases on some tasks but drops on others., indicating that parameter initialisation is a key to success in knowledge distillation. The last three columns in Table 6 show the average scores on BUCC and Tatoeba for each language pair, indicating that PCA and contrastive learning in Section 3.3 are very effective on cross-lingual tasks. It also reveals that ABTT, i.e. removing the first $r$ PCs from SWEs, plays a vital role in generating cross-lingual SWEs.
Table 7 shows performance of our models trained with different STs, namely SimCSE, Nomic Nussbaum et al. (2024), and GTE-{small/base/large}. It indicates that OURS performs better using better STs (SimCSE < Nomic < GTE) as its teacher model, suggesting that performance of our model might keep increasing as ST models improve. That being said, our model does not benefit from using the largest GTE model, possibly due to the large discrepancy of the parameter sizes between the teacher and student models during knowledge distillation. Performance on each MTEB task is shown in Table 12 in Appendix.
5.2 Effects of PCA
$i$ Smallest Largest 1 advantages, maps, permits, categories, manufacturers haha, ha, grinned, omg, oh 2 appreciate, hey, haha, btw, anyways shook, sheriff, snapped, glanced, sighed 6 nov, oct, dec, 2020, feb charming, compliment, courteous, friendly, softly
Table 8: Examples of words with the smallest and largest values in the $i$-th principal component.
To investigate why PCA with ABTT is particularly effective on cross-lingual tasks, we analyse what information is encoded in the first two PCs and find that it captures the variance largely based on the language identity. To visualise this, we randomly sample 1,000 language-specific words/tokens from English and Japanese and plot their embeddings based on the values in the first two PCs (PC1 and PC2). The result is presented in Figure 1 (the leftmost plot), which clearly shows that the embeddings are separated based on language. Given this finding, we also visualise the embeddings using t-SNE van der Maaten and Hinton (2008) before and after applying PCA wo/w ABTT, and the results are also shown in Figure 1. We can see that the embeddings create language clusters before PCA and after PCA without ABTT, but applying PCA with ABTT renders the embeddings more language-agnostic by removing the language-identity components. We also observe the same trend in English–Chinese/German, as shown in Figures 5 and 6 in Appendix. Such language-specific dimensions can act as noise when measuring semantic similarity across different languages, [^9] For instance, Japanese sentences with many English loanwords can have high similarities to English sentences regardless of the semantics. explaining why ABTT is crucial in cross-lingual SWEs. Furthermore, this observation makes us wonder whether such language segregation also occurs in the mGTE embedding space; we investigate this and find that it does create language clusters at lower layers of Transformer, but generates more language-agnostic embeddings at higher layers. Since it deviates from the main subject of this study, we discuss more details in Appendix A.
Lastly, we also analyse the dominant PCs of monolingual (English) SWEs, and find that some of them seem to capture linguistic styles or domains. Table 8 shows examples of the words that have the five largest and smallest values in PC1, PC2, and PC6 among the 10k most frequent words in $V$. In PC1 and PC2, words seem to be clustered based on the word frequency in casual text, whereas in PC6, date-related words seem to have large values. To verify whether PC1 and PC2 are actually related to textual formality, we conduct a simple experiment using the data set released by Chhaya et al. (2018), where 10 human annotators judge the formality of 960 emails. We encode each email by averaging SWEs and measure the Pearson correlation coefficient between the values in PC1, PC2, or PC6, and the mean scores of the 10 annotators. We observe that the correlations are: $-0.63$, $0.37$, and $-0.11$ for PC1, PC2, and PC6, confirming a moderate or weak correlation for the first two PCs. [^10] We surmise that this may be attributed to the use of sentences sampled from a web corpus in our model, which is a mixed bag of formal and casual text. It would be interesting to see how the selection of text data affects the outcome. Albeit those components capture the variance of sentences effectively, they would not necessarily serve as useful features on semantic tasks, explaining why PCA w/ ABTT works well for monolingual SWEs as well, [^11] Yet, ABTT often harms performance on classification. which Mu and Viswanath (2018) describe as counter-intuitive.
5.3 Analysis of Word Embedding Norm
Figure 2: Comparison of word embedding norms obtained in Sections 3.1, 3.2, and 3.3. For normalisation, norms of each POS tag are divided by the maximum value of all tags (i.e. the norm of Proper Noun).
To investigate how our method optimises word embeddings, we analyse the norms of SWEs obtained in Sections 3.1, 3.2, and 3.3; we denote them as $\|\mathrm{E}_{i}(w)\|$ ($i=1,2,3$). For the purpose of analysis, we compare the norms of the 10k most frequent words in Brown Corpus, which consists of text in various domains and provides part-of-speech (POS) tags for each word. First, we compare the mean norms of word embeddings for each POS tag, [^12] We use the most frequent tag for each word. and Figure 2 shows the results. For normalisation, we divide the norms of each POS tag by that of “Proper Noun”, which always has the largest value regardless of $i$. It shows that the differences in the norms are emphasised after PCA and knowledge distillation. In particular, the relative norms of “Article” (e.g. the), “Preposition” (e.g. of), and “Numeral” (e.g. 2025, first) drop sharply because these types of words do not affect sentence semantics very much. On the other hand, “Proper Noun” and “Noun” have the largest norms after PCA and knowledge distillation, as they often convey important information. This result is somewhat akin to the previous finding that ST tends to emphasise nominal information of a sentence Nikolaev and Padó (2023). Lastly, we find that our method induces a similar effect to SIF, a heuristic used in Model2Vec — we measure Spearman’s rank correlation between the norm and word frequency rank and find that it increases after PCA and knowledge distillation ($0.39\rightarrow 0.43\rightarrow 0.49$), meaning both steps decrease the norms of high-frequency words. Compared to SIF, however, knowledge distillation adjusts norms in a more nuanced way, considering not only word frequency (which is implicitly encoded by ST Kurita et al. (2023)) but also POS tags and influence of each word on sentence semantics.
5.4 Impact of Model Size
$d$ $|V|$ Avg-s2s Avg 512 150k 64.1 52.4 256 150k 63.8 52.0 128 150k 62.6 50.4 64 150k 52.5 41.4 256 100k 63.5 51.8 256 75k 63.4 51.5 256 50k 62.7 51.0 256 30k 62.2 50.2
Table 9: Performance of our model on MTEB with different embedding and vocabulary sizes.
Table 9 compares OURS trained with different dimensionalities $d$ and vocabulary sizes $|V|$. Our model performs the best with $d\geq 256$ and also achieves good scores with $d=128$. However, the scores drop sharply with $d=64$, where knowledge distillation rather harms performance (44.0 $\rightarrow$ 41.4) likely due to the large discrepancy of the embedding spaces between the teacher and student models. In terms of $|V|$, our model performs well with $|V|\geq 75\mathrm{k}$. Hence, setting $d=256$ and $|V|\geq 75\mathrm{k}$ would offer a good balance between performance and memory usage (in English). We provide scores on each task in Table 13 in Appendix.
6 Conclusion
We propose new word embeddings optimised for sentence semantic representation. We extract embeddings from Sentence Transformers and improve them with sentence-level PCA and knowledge distillation or contrastive learning. Our rigorous experiments show that our model outperforms baselines on both monolingual and cross-lingual tasks. Our analyses also reveal that our method removes embedding components that are not relevant to sentence semantics, and adjusts the vector norms based on the influence on sentence semantics.
7 Limitations
Since our model (and other SWEs) does not consider word order information (i.e. a bag-of-words model), its performance is limited on tasks where models are required to understand long context. Importantly, however, there is always a trade-off between model performance and computational complexity; in this work, we focus on improving the cost-efficiency, an important factor for real-world applications that is nonetheless largely neglected among the research community these days. To improve performance (at the cost of computational cost), one can simply add a small network that can capture sequential information (e.g. Transformer, CNN, LSTM Hochreiter and Schmidhuber (1997)) on top of our word embeddings, and train them with labelled data following the training process of Sentence-Transformer models. Another idea to better capture sentence semantics is to represent text with a set of word embeddings with positional information (rather than taking their average), and calculate textual similarity by measuring the distance (or optimal transport cost) between two distributions, e.g. Word Mover’s Distance Kusner et al. (2015).
Appendix A Analysis of mGTE Embeddings
Figure 3 shows the t-SNE visualisation of sentence embeddings produced by mGTE-base at the 1st, 4th, 8th, and 12th (= last) layers of Transformer. We encode 1,000 translation pairs sampled from an English–Japanese parallel corpus; as such, ideally they should have close representations regardless of language. To obtain sentence vectors, we apply mean pooling at each layer. [^13] For the last layer, we also tried using the [[[[CLS]]]] embedding following the default pooling configuration, and got a similar result. The figure shows that the embeddings are completely separated by language at the 1st and 4th layers (just like our SWEs before PCA in Figure 1), but are less clustered at the 8th layer and completely mixed at the the last layer. [^14] As such, applying PCA did not improve mGTE’s performance. This suggests that mGTE recognises the input language at lower layers, and then captures language-agnostic features such as sentence semantics at higher layers. We also verify this hypothesis by evaluating embeddings at each layer on a translation retrieval task and observing better performance at higher layers.
Notably, this result is somewhat inconsistent with the previous findings that multilingual masked language models (MLMs) like mBERT have language-specific subspaces in their embedding even at higher layers Gonen et al. (2020); Libovický et al. (2020). One possible reason for this is that mGTE is fine-tuned on various tasks to produce sentence representation, which might make it generate language-agnostic embeddings. To verify this hypothesis, we also evaluate the embeddings of mGTE’s backbone MLM (i.e. a BERT-like model with several enhancements) before it was fine-tuned on labelled data, [^15] We use the model at https://huggingface.co/Alibaba-NLP/gte-multilingual-mlm-base. and Figure 4 shows the results. It clearly indicates that the embeddings are completely separated based on language at all layers, suggesting that mGTE benefits from fine-tuning in producing language-agnostic representations. Before fine-tuning, multilingual MLMs are pre-trained to predict masked tokens in each language at the last layer, and that would naturally encourage them to have language-specific subspaces, as discussed in Gonen et al. (2020).
Models Emb Size $d$ Model Path MiniLM-L6 384 sentence-transformers/all-MiniLM-L6-v2 GTE-small 384 thenlper/gte-small GTE-base 768 Alibaba-NLP/gte-base-en-v1.5 GTE-large 1,024 Alibaba-NLP/gte-large-en-v1.5 mGTE 768 Alibaba-NLP/gte-multilingual-base mGTE-MLM 768 Alibaba-NLP/gte-multilingual-mlm-base BGE 768 BAAI/bge-base-en-v1.5 SimCSE 768 princeton-nlp/sup-simcse-bert-base-uncased Nomic 768 nomic-ai/nomic-embed-text-v1.5 LLaMa3 4,096 McGill-NLP/LLM2Vec-Meta-Llama-3-8B-Instruct-mntp-supervised fastText 300 crawl-300d-2M.vec Sent2vec 700 sent2vec_wiki_bigrams WordLlama 512 wordllama-l3-supercat Model2Vec 512 potion-base-32M Model2Vec (cross-lingual) 256 M2V_multilingual_output
Table 11: Details of the models used in our experiments.
Models ($d$) Class. Clust. PairClass. Rerank. Retr. STS Summ. Avg-s2s Avg # Data Sets 12 11 3 4 15 10 1 33 56 Sentence Transformers (STs) SimCSE-supervised (768) 67.32 33.43 73.68 47.54 21.82 79.12 31.17 62.75 48.86 Nomic (768) 73.55 43.93 84.61 55.78 53.01 81.94 30.40 69.52 62.28 GTE-small (384) 72.31 44.89 83.54 57.70 49.46 82.07 30.42 69.32 61.36 GTE-base (768) 77.17 46.82 85.33 57.66 54.09 81.97 31.17 71.51 64.11 GTE-large (1,024) 77.75 47.96 84.53 58.50 57.91 81.43 30.91 71.65 65.39 Static Word Embeddings (SWEs) OURS-SimCSE (256) 64.72 32.78 75.65 48.36 26.54 74.83 30.61 61.16 48.83 OURS-Nomic (256) 65.94 34.74 79.29 50.15 30.47 75.49 31.49 62.69 50.99 OURS-GTE-small (256) 63.90 34.14 78.17 49.66 27.39 72.63 31.14 60.76 49.00 OURS-GTE-base (256) 67.29 36.67 78.99 50.91 31.32 75.87 30.55 63.76 51.97 OURS-GTE-large (256) 65.36 36.11 78.13 50.42 31.58 75.59 30.73 62.91 51.39
Table 12: Performance of different ST models and our models ($d=256$) distilled from each ST.
Models ($d$) Class. Clust. PairClass. Rerank. Retr. STS Summ. Avg-s2s Avg # Data Sets 12 11 3 4 15 10 1 33 56 OURS ($d$ = 64,   $|V|$ = 150k) 60.20 29.09 60.12 44.07 17.83 62.12 30.07 52.51 41.39 OURS ($d$ = 128, $|V|$ = 150k) 65.92 35.51 78.05 50.09 28.33 75.06 30.60 62.62 50.40 OURS ($d$ = 256, $|V|$ = 150k) 67.29 36.67 78.99 50.91 31.32 75.87 30.55 63.76 51.97 OURS ($d$ = 512, $|V|$ = 150k) 67.96 37.14 79.00 51.07 32.01 75.59 30.35 64.06 52.35 OURS ($d$ = 256, $|V|$ = 150k) 67.29 36.67 78.99 50.91 31.32 75.87 30.55 63.76 51.97 OURS ($d$ = 256, $|V|$ = 100k) 67.29 36.12 78.56 50.71 31.18 75.90 30.84 63.55 51.80 OURS ($d$ = 256, $|V|$ = 75k) 67.38 35.67 78.55 50.59 30.56 75.67 30.36 63.36 51.51 OURS ($d$ = 256, $|V|$ = 50k) 67.26 34.11 78.19 50.30 30.25 75.21 30.68 62.70 50.97 OURS ($d$ = 256, $|V|$ = 30k) 66.90 32.89 78.03 50.00 28.61 75.14 29.81 62.25 50.16
Table 13: Performance of OURS trained with different embedding and vocabulary sizes.
Appendix B Word Embedding Ensembling
Given the success of ensemble learning on many NLP tasks, we try combining $L$ SWEs trained with different algorithms. For the input text $z$, we generate its ensemble embedding $\mathrm{f}(z)$ as:
$\mathrm{f}(z)$ $=\frac{1}{\sqrt{\sum_{i=1}^{L}\lambda_{i}}}\left[\frac{\sqrt{\lambda_{i}}}{\|\mathrm{f}_{i}(z)\|}\mathrm{f}_{i}(z)\right]_{i=1}^{L}~{},$ (6)
$\mathrm{f}_{i}(z)$ $=\frac{1}{|z|}\sum_{w^{\prime}\in z}\mathrm{E}_{i}(w^{\prime}),$ (7)
[…]
The last three rows in Table 10 show the results when we ensemble OURS with WordLlama (WL) and Model2Vec (M2V); regarding the last row 2*OURS + WL + MV, we set $\lambda_{i}$ to $2$ for OURS and $1$ for the other models to increase the contribution of OURS, the best-performing SWE. It shows that it performs the best of all models in both Avg-s2s and Avg, indicating the effectiveness of our ensembling method. Looking at the scores on each task, our ensembling method successfully combines the strengths of each model, e.g. the ensembled model substantially improves performance over OURS on Reranking and Retrieval, where WordLlama performs the best of all SWEs.
Appendix C Model Details
Table 11 lists the details of the models used in our paper. The Sent2vec model is downloaded from https://github.com/epfml/sent2vec; fastText from https://fasttext.cc/docs/en/english-vectors.html Mikolov et al. (2018); and the rest from Hugging Face (https://huggingface.co/models).